# Quantum Doubly Stochastic Transformers

### v1

@@ 2.3.2 Projection on the Birkhoff polytope
Alternatively, one can project $\bm{\mathrm{M}}$ directly on $\Omega_{T}$ via $\bm{\mathrm{P}}=\underset{\bm{\mathrm{X}}\in\Omega_{T}}{\operatorname*{arg\,min}}\|\bm{\mathrm{X}}-\bm{\mathrm{M}}\|_{F}^{2}$, where the set for $\bm{\mathrm{X}}$ and the objective are convex. We chose to minimize the Frobenius norm here but note that different distances could be explored. The resulting problem is a positive-definite convex quadratic program and can be rewritten as
$\min_{{\begin{subarray}{c}s.t.\\ \bm{\mathrm{x}}_{i}\geq 0\\ \bm{\mathrm{A}}\bm{\mathrm{x}}=\mathbf{1}_{2n}\end{subarray}}}\tfrac{1}{2}\bm{\mathrm{x}}^{\top}\bm{\mathrm{x}}-\bm{\mathrm{q}}^{\top}\bm{\mathrm{x}},\quad\bm{\mathrm{A}}={\text{\scriptsize$\begin{pmatrix}\bm{\mathrm{1}}_{n}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{0}}^{\top}\\ \bm{\mathrm{0}}_{n}^{\top}&\bm{\mathrm{1}}_{n}^{\top}&\ldots&\bm{\mathrm{0}}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{\mathrm{0}}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{1}}_{n}^{\top}\\ \bm{\mathrm{I}}_{n}&\bm{\mathrm{I}}_{n}&\ldots&\bm{\mathrm{I}}_{n}\\ \end{pmatrix}$}}$ (5)
[…]
@@ 2.3.3 QontOT
QontOT is a parameterized (variational) quantum circuit that was conceived very recently for conditional prediction of optimal transport plans [37]. Importantly, the circuit naturally emits DSMs which are rescaled with classical methods to transport plans. While [37] do not find signs of quantum advantage for their main task of optimal transport plan prediction, they report accuracy surpassing their classical baselines for the prediction of DSMs. This is likely a consequence of the choice of the ansatz which explores a previously unreported link between unitary operators and DSMs. Indeed, they first proved that DSMs can be obtained naturally with quantum computers thus constructing a quantum inductive bias for DSMs. Notably, as the authors state, it is currently unknown whether a similarly natural classical approach exists to produce DSMs parametrically.
For any unitary matrix $\bm{\mathrm{U}}$: $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$. Given the circuit parameters $\theta$ (typically in the hundreds) and $p\in\mathbb{R}$, QontOT obtains a DSM via $\bm{\mathrm{U}}(p;\theta)\odot\overline{\bm{\mathrm{U}}}(p;\theta)$. Mariella et al. [37] block-decompose this matrix before the classical rescaling. A notable detriment of QontOT is the data injection which is limited to a scalar $p$. Therefore, we extend the multiplicative data injection $f(\theta,p)=p\cdot\theta$ from scalars to tensors, such that $f(\theta,\bm{\mathrm{M}})=\theta\odot\overrightarrow{\bm{\mathrm{M}}}$. If $\bm{\mathrm{M}}$ has less items than $\theta$ we repeat its values to obtain a vector of length identical to $\theta$. Furthermore, QontOT requires the DSM dimension $n$ to be a power of $2$. While this may be prohibitive within a Transformer (because sequence length $T$ may differ), it can be mitigated by padding. Padding to powers of two is a common technique to maximize hardware efficiency. Here, we focus our experiments on ViTs where $T$ is a direct function of patch size. In general, the circuit size scales favorably in $\mathcal{O}(\log_{2}(T))$, to be precise, it needs at least $4(\log_{2}(T)+1)$ qubits, equivalent to $2(q_{d}+q_{a})$ where $q_{d}$ is the number of data qubits ($\log_{2}(T)$) and $q_{a}$ is the number of auxilliary qubits ($\geq\log_{2}(T)+1$). In practice $T=8$.
@@ 2.3.4 QR Decomposition
As highlighted above, any unitary $\bm{\mathrm{U}}$ can provide a DSM by taking $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}$. For any input matrix $\bm{\mathrm{M}}$, we can obtain a unitary $\bm{\mathrm{U}}$ by computing an orthonormal basis for its column space. While there are many ways to obtain a basis, we choose a QR decomposition $\bm{\mathrm{M}}=\bm{\mathrm{U}}\bm{\mathrm{R}}$, in which case $\bm{\mathrm{R}}$ is upper triangular. [^1] We use $\bm{\mathrm{U}}$ rather than $\bm{\mathrm{Q}}$ for the basis of the QR to avoid confusion with the Transformer query matrix. QR has the advantage that, when implemented with Gram-Schmidt, it is differentiable if $\bm{\mathrm{M}}$ is full-rank. However note that for long-context applications the rank is rarely full because the query and key matrix have $d=\frac{d_{embed}}{n_{heads}}$ rows and typically $\bm{\mathrm{M}}\in\mathbb{R}^{T\times T}$ has rank $\min\{d,T\}$, implying that $\bm{\mathrm{M}}$ only has full rank when $d\geq T$. In practice, if the rank is defective, we inject additive Gaussian noise $\mathcal{N}(0,1\mathrm{e}{-7})$ to obtain full ranks. In the ViTs studied in our experiments, $\bm{\mathrm{M}}$ often has close to, or full-rank since the dimension $d$ is greater than the number of patches $P$, where $P\approx T$. Moreover, QR has time complexity $\mathcal{O}(n^{3})$ for dense $n\times n$ matrices, thus to scale up, approximation techniques may be needed [22].
@@ 3 Expressivity of Doubly-Stochastic Operators
Given the empirical superiority of the Sinkformer to the vanilla (i.e., right-stochastic) Transformers, a natural question is which operator to chose to obtain DSMs. Before training the DSFormers, we compare the expressivity of the operators – especially QontOT and Sinkhorn’s algorithm – in isolation on synthetic data. We focus on two aspects.
@@
    Soundness – does the operator always produce a DSM? Given that $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$, QontOT always yields a DSM. Similarly for the QR decomposition. Instead, Sinkhorn’s algorithm (SA) may fail to produce a DSM if the input matrix is not positive. Within the Transformer where $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$, the positivity requirement is generally not fulfilled which is mitigated by exponentiation of the inputs. Thus, following Sinkhorn’s theorem, SA always converges given enough iterations $k$. In practice however, the iterative procedure is limiting. When passing $8\times 8$ $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ matrices from a trained Sinkformer, we observe that SA does not converge for the common choices of $k$ ($3$ and $21$). Indeed, the Frobenius distance to the Birkhoff polytope is $0.84_{\pm 0.26}$ for $k=3$ and $0.23_{\pm 0.16}$ for $k=21$. This is in contrast to the QR, QontOT and the Birkhoff projection which all yield distances $<2\mathrm{e}{-4}$, QontOT even $<5\mathrm{e}{-6}$. Instead, the vanilla Softmax operator yields a right-stochastic matrix with distance $1.12_{\pm 0.33}$. Hence Sinkformer attention is only approximately doubly stochastic.
@@
    Completeness – can the operator produce all possible DSMs? Sinkhorn’s algorithm reaches all DSMs of the form $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}$. However, due to the entry-wise exponentiation of $\bm{\mathrm{M}}$ in the Sinkformer, the input matrix never contains any zero, thus the boundaries of the polytope cannot be reached. Regarding QontOT, the resulting DSM is a convex combination of unistochastic matrices [37][Eq. 11b]. Unistochastic matrices are a non-convex proper subset, covering a large amount of the Birkhoff polytope (albeit the exact amount is unknown [18]). In theory, if all unistochastic matrices could be reached, then by their convex combinations QontOT could cover the entire Birkhoff polyope. In practice, reaching all unistochastic matrices (especially all permutation matrices) with the same circuit parametrization is unfeasible as it requires fault-tolerant quantum hardware and high circuit depth (entanglement). But over the parameter space of QontOT, the Birkhoff polytope can be approximated more closely.
@@ 3.1 Empirical analysis
Figure 2: Number of unique DSMs obtained with each operator after exhaustively iterating over a discretized unit hypercube. With only $8$ layers, QontOT produces a unique DSM for every possible input, unlike all other methods.
@@
To empirically assess the completeness of the operators w.r.t. the Birkhoff polytope, we performed a brute-force analysis over a discretized grid of the unit hypercube. For a $n\times n$ matrix and a discretiztion step $d\in\mathbb{N}_{+}$, we sample each column from a discretized $n$-dimensional hybercube with $d^{n}$ points, yielding $d^{n^{2}}$ unique matrices. For $n=4$ and $d=3$ we obtain $3^{16}\approx 43$M matrices and computed the DSM for each input, before rounding to third decimal place. Across all operators, QontOT yielded by far the most unique DSMs (see Figure 2), behaving nearly injective when $8$ or more circuit layers are used. This is important, because, none of the other operators is injective thus some information is lost when using it inside a neural network. A closer inspection of the empirical cumulative distribution function of all DSMs reveals that QR often emits the same DSMs and that with only $2$ layers (i.e., $98$ parameters), QontOT surpasses all other methods (see Appendix Figure A2). Furthermore, whereas all the classical techniques are non-parametric, QontOT is a learning method, thus each parameter configuration yields a different set of DSMs. In general, Sinkhorn’s algorithm and the direct Birkhoff projection are both permutation and rotation equivariant. Instead, QR and QontOT do not possesses this characteristic. Moreover, the Naive flavor of SA and the QR are scale-invariant whereas OT and QontOT are not. Note that such equivariances within a Transformer are not generally beneficial or detrimental. We also examined the operators on $8\times 8$ $\bm{\mathrm{Q}}\bm{\mathrm{K}}$ matrices from a single batch ($N=3584$, $7$ heads, batch size of $512$) of a trained Sinkformer. In general, the operators were injective on this set, yielding unique DSMs, however the OT flavor of Sinkhorn’s algorithm only produced $2175$ ($2021$) unique DSMs for $k=3$ ($21$) iterations. This deficiency may be due to the inputs having large numerical values and not being generally non-negative.
Beyond approximate injectivity, we argue that a powerful operator needs to possess two further characteristics: First, information has to be preserved. Obtaining unique DSMs is useless if they destroy information from the input matrix. To assess this, we measured the Frobenius norm of the residuals between input and output matrix [^2] Other metrics like measuring preservation of ranks or pairwise ratios are possible but yielded similar results.. Secondly, low entropy has to be avoided because it causes vanishing gradients and destabilizes Transformer training [62, 45] for which various mitigation techniques have been suggested [55, 24]. This so-called “entropy collapse” arises if attention is too spiky and is induced by low temperature in the softmax. Our analysis in Figure 3 reveals that QontOT possess better information preservation and larger entropy than Sinkhorn’s algorithm on realistic unnormalized attention matrices.
@@
Figure 3: Shannon entropy of different flavors of doubly stochastic attention against a measure of distance preservation – the Frobenius norm of the difference between unnormalized attention $\bm{\mathrm{Q}}\bm{\mathrm{K}}$ and their obtained DSM $\bm{\mathrm{P}}$. QontOT uses $2$ circuit layers.
Among all operators, QontOT possess the best information preservation properties while still maintaining Shannon entropy that is only surpassed by the QR decomposition.
@@
Altogether, these results are promising and motivate our below empirical experiments with a Quantum Doubly Stochastic Transformer (QDSFormer).

### v2

@@ 2.3.2 Projection on the Birkhoff polytope
Previous work studied different approaches to project matrices onto the Birkhoff polytope [lim2020doubly, ding2022understanding], and the most established scheme leverages Frobenius distance [zass2006doubly]. Alternatively, one can project $\bm{\mathrm{M}}$ directly on $\Omega_{T}$ via $\bm{\mathrm{P}}=\underset{\bm{\mathrm{X}}\in\Omega_{T}}{\operatorname*{arg\,min}}\|\bm{\mathrm{X}}-\bm{\mathrm{M}}\|_{F}^{2}$, where the set for $\bm{\mathrm{X}}$ and the objective are convex. We chose to minimize the Frobenius norm here but note that different distances could be explored. The resulting problem is a positive-definite convex quadratic program and can be rewritten as
$\min_{{\begin{subarray}{c}s.t.\\ \bm{\mathrm{x}}_{i}\geq 0\\ \bm{\mathrm{A}}\bm{\mathrm{x}}=\mathbf{1}_{2n}\end{subarray}}}\tfrac{1}{2}\bm{\mathrm{x}}^{\top}\bm{\mathrm{x}}-\bm{\mathrm{q}}^{\top}\bm{\mathrm{x}},\quad\bm{\mathrm{A}}={\text{\scriptsize$\begin{pmatrix}\bm{\mathrm{1}}_{n}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{0}}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{\mathrm{0}}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{1}}_{n}^{\top}\\ \bm{\mathrm{I}}_{n}&\bm{\mathrm{I}}_{n}&\ldots&\bm{\mathrm{I}}_{n}\\ \end{pmatrix}$}}$ (3)
where $\bm{\mathrm{x}}=\mathrm{vec}(\bm{\mathrm{X}}^{T})$, $\bm{\mathrm{q}}=\mathrm{vec}(\bm{\mathrm{M}}^{T})$ and $\bm{\mathrm{A}}\in\mathbb{R}^{2n\times n^{2}}$. The last (or any other) row of $\bm{\mathrm{A}}$ can be removed without losing information, since it is a linear combination of the other rows ($\bm{\mathrm{A}}$ has rank 2n1)2n-1) [brualdi_2006, Thm. 8.1.1]. We solved the quadratic program with OSQP [stellato2020osqp].
@@ 2.3.3 QontOT
QontOT is a parameterized (variational) quantum circuit that was conceived for conditional prediction of optimal transport plans [mariella2024quantum] but can be extended to many combinatorial problems [mermoud2025variational]. The circuit naturally emits DSMs and while [mariella2024quantum] do not find signs of quantum advantage for their main task of optimal transport plan prediction, they report accuracy surpassing their classical baselines for the prediction of DSMs. This is likely a consequence of the choice of the ansatz which explores a previously unreported link between unitary operators and DSMs. Indeed, they first proved that DSMs can be obtained naturally with quantum computers thus constructing a quantum inductive bias for DSMs. Notably, as the authors state, it is currently unknown whether a similarly natural classical approach exists to produce DSMs parametrically.
Let $\odot$ be the Hadamard product and $\overline{U}=\left(U^{\dagger}\right)^{\top}$ the complex conjugate. For any unitary matrix $\bm{\mathrm{U}}$: $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$. Given the circuit parameters $\theta$ (typically in the hundreds) and $p\in\mathbb{R}$, QontOT obtains a DSM via $\bm{\mathrm{U}}(p;\theta)\odot\overline{\bm{\mathrm{U}}}(p;\theta)$. This matrix is block-decomposed before the classical rescaling. A notable detriment of QontOT is the data injection which is limited to a scalar $p$. Therefore, we extend the multiplicative data injection $f(\theta,p)=p\cdot\theta$ from scalars to tensors, such that $f(\theta,\bm{\mathrm{M}})=\theta\odot\overrightarrow{\bm{\mathrm{M}}}$. If $\bm{\mathrm{M}}$ has less items than $\theta$ we repeat its values to obtain a vector of length identical to $\theta$. Furthermore, QontOT requires the DSM dimension $n$ to be a power of $2$. While this may be prohibitive within a Transformer (because sequence length $T$ may differ), it can be mitigated by padding. Padding to powers of two is a common technique to maximize hardware efficiency. Here, we focus our experiments on ViTs because $T$ is a function of patch size. In general, the circuit size scales favorably in $\mathcal{O}(\log_{2}(T))$. It needs at least $4(\log_{2}(T)+1)$ qubits, i.e., $2(q_{d}+q_{a}+1)$ where $q_{d}$ is the number of data qubits ($\log_{2}(T)$) and $q_{a}$ is the number of auxilliary qubits ($\geq\log_{2}(T)+1$).
@@ 2.3.4 QR Decomposition
As highlighted above, any unitary $\bm{\mathrm{U}}$ can provide a DSM by taking $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}$. For any input matrix $\bm{\mathrm{M}}$, we can obtain a unitary $\bm{\mathrm{U}}$ by computing an orthonormal basis for its column space. While there are many ways to obtain a basis, we choose a QR decomposition $\bm{\mathrm{M}}=\bm{\mathrm{U}}\bm{\mathrm{R}}$, in which case $\bm{\mathrm{R}}$ is upper triangular. When implemented with Gram-Schmidt, QR is differentiable if $\bm{\mathrm{M}}$ is full-rank, but for long-context applications that is rarely the case because query and key matrix have $d=\frac{d_{embed}}{n_{heads}}$ rows and typically $\bm{\mathrm{M}}\in\mathbb{R}^{T\times T}$ has rank $\min\{d,T\}$, implying that $\bm{\mathrm{M}}$ only has full rank when $d\geq T$. In practice, if the rank is defective, we inject additive Gaussian noise $\mathcal{N}(0,1\mathrm{e}{-7})$ to obtain full ranks. In the ViTs studied in our experiments, $\bm{\mathrm{M}}$ often has close to, or full-rank since the dimension $d$ is greater than the number of patches $P$, where $P\approx T$. Moreover, QR has time complexity $\mathcal{O}(n^{3})$ for dense $n\times n$ matrices, thus to scale up, approximation techniques may be needed [halko2011finding].
@@ 3 Expressivity of Doubly-Stochastic Operators
Given the empirical superiority of the Sinkformer to the vanilla (i.e., right-stochastic) Transformer, a natural question is which operator to chose to obtain DSMs. Before training the DSFormers, we compare the expressivity of the operators – especially QontOT and Sinkhorn’s algorithm – in isolation on synthetic data. We focus on two aspects.
@@
    Soundness – does the operator always produce a DSM? Given that $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$, QontOT always yields a DSM. Similarly for the QR decomposition. Instead, Sinkhorn’s algorithm (SA) may fail to produce a DSM if the input matrix is not positive. Within the Transformer where $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$, the positivity requirement is generally not fulfilled which is mitigated by input exponentiation. Thus, following Sinkhorn’s theorem, SA always converges given enough iterations $k$. But in practice the iterative procedure is limiting. When passing $8\times 8$ $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ matrices from a trained Sinkformer, we observe that SA does not converge for the common choices of $k$ (33 and $21$). Indeed, the Frobenius distance to the Birkhoff polytope is $0.84_{\pm 0.3}$ for $k=3$ and $0.23_{\pm 0.2}$ for $k=21$. This is in contrast to the QR, QontOT and the Birkhoff projection which all yield distances $2\mathrm{e}{-4}$, QontOT even $5\mathrm{e}{-6}$. Instead, the vanilla Softmax operator yields a right-stochastic matrix with distance $1.12_{\pm 0.3}$. Hence Sinkformer attention is only approximately doubly stochastic.
@@
    Completeness – can the operator produce all possible DSMs? Sinkhorn’s algorithm reaches all DSMs of the form $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}$. However, due to the entry-wise exponentiation of $\bm{\mathrm{M}}$ in the Sinkformer, the input matrix never contains any zero, thus the boundaries of the polytope cannot be reached. Regarding QontOT, the resulting DSM is a convex combination of unistochastic matrices [mariella2024quantum, Eq. 11b]. Unistochastic matrices are a non-convex proper subset, covering a large amount of the Birkhoff polytope (albeit the exact amount is unknown [dunkl2009volume]). In theory, if all unistochastic matrices could be reached, then by their convex combinations QontOT could cover the entire Birkhoff polyope. In practice, reaching all unistochastic matrices (especially all permutation matrices) with the same circuit parametrization is unfeasible as it requires fault-tolerant quantum hardware and high circuit depth (entanglement). But over the parameter space of QontOT, the Birkhoff polytope can be approximated more closely.
@@ 3.1 Empirical analysis
To empirically assess the completeness of the operators w.r.t. the Birkhoff polytope, we performed a brute-force analysis over a discretized grid of the unit hypercube. For a $n\times n$ matrix and a discretiztion step $d\in\mathbb{N}_{+}$, we sample each column from a discretized $n$-dimensional hybercube with $d^{n}$ points, yielding $d^{n^{2}}$ unique matrices. We refrain from analysing vectors above unit length because all operators are scale-invariant, i.e., $f(\lambda A)=f(A)$. For $n=4$ and $d=3$ we obtain $3^{16}\approx 43$M matrices and computed the DSM for each input, before rounding to third decimal place. Across all operators, QontOT yielded by far the most unique DSMs (see Figure 2A), behaving nearly injective when $8$ or more circuit layers are used. This is important, because, none of the other operators is injective thus some information is lost when using it inside a neural network. A closer inspection of the empirical cumulative distribution function of all DSMs reveals that QR often emits the same DSMs and that with only $2$ layers (i.e., $98$ parameters), QontOT surpasses all other methods (see Appendix Figure A3). Furthermore, whereas all the classical techniques are non-parametric, QontOT yields a different set of DSMs for each parameter configuration. We repeated above experiment by sampling from a discretized grid of the unit hypersphere (instead of the hypercube). In this case, Sinkhorn also produces collisions, while QontOT remains injective (there is no proof that the map is injective, but it appears to be experimentaly). With a discretization of $d=3$, our sphere contains $625$ matrices where all columns have unit lengths. Sinkhorn yields collisions by mapping all rank-1 matrices with constant rows to the center of the Birkhoff polytope, i.e., it fails to differentiate matrix $\mathbf{e}_{2}\mathbf{1}^{\top}$ and $\mathbf{e}_{4}\mathbf{1}^{\top}$ ($\mathbf{e}_{2}$ and $\mathbf{e}_{4}$ are the second and forth column of the identity) and thus produces only $621$ unique DSMs whereas QontOT yields $625$ DSMs (QR: 381). This is critical because it implies that Sinkhorn confuses cases where attention matrices are row-wise constant but each row has a unique value. In general, Sinkhorn’s algorithm and the direct Birkhoff projection are both permutation and rotation equivariant. Instead, QR and QontOT do not possess this characteristic. Moreover, SA and the QR are scale-invariant whereas, again, QontOT is not. Note that such invariances or equivariances within a Transformer are not generally beneficial or detrimental.
Beyond approximate injectivity, a powerful operator needs to possess two further characteristics: First, information has to be preserved. Obtaining unique DSMs is useless if they destroy information from the input matrix.
@@
Figure 2: Left: Number of unique DSMs obtained after exhaustively iterating over a discretized unit hypercube. With only 88 layers, QontOT produces a unique DSM for every possible input, unlike all other methods. Right: Entropy vs. distance-preservation tradeoff. Shannon entropy of different doubly stochastic attention against the Frobenius norm of the difference between unnormalized attention $\bm{\mathrm{Q}}\bm{\mathrm{K}}$ and the obtained DSM $\bm{\mathrm{P}}$.
To assess this, we measured the Frobenius norm of the residuals between input and output matrix [^1] Other metrics like measuring preservation of ranks or pairwise ratios are possible but yielded similar results.. Secondly, low entropy has to be avoided because it causes vanishing gradients and destabilizes Transformer training [zhai2023stabilizing, shen2023study] for which various mitigation techniques have been suggested [wang2021escaping, hoffmann2024eureka]. This so-called “entropy collapse” arises if attention is too spiky and is induced by low temperature in the softmax. Our analysis in Figure 2B reveals that QontOT possesses comparable information preservation to Sinkhorn while having higher entropy on realistic unnormalized attention matrices. QR decomposition showed superior entropy, but its cubic scaling limits applicability beyond small-scale Transformers.